OpenStudy (anonymous):

The first challenge to complete is the Penny Problem. The radio station is giving the winner of this challenge a prize pack that includes tickets to see his or her favorite band in concert. To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? You must show all work to receive credit.

OpenStudy (anonymous):

penny diameter = 0.75 radius = ? 0.375 height = 0.061 base = (3.14)(0.375)^2 = 0.44 volume = (0.44)(0.061) = 0.027 glass jar diameter = 6 radius = ? 3 height = 11.5 base = (3.14)(3)^2 = 28.26 volume = (28.26)(11.5) = 324.99 | 325 am i right?

OpenStudy (anonymous):

truth

OpenStudy (anonymous):

@shelbygt520 @iPwnBunnies @lovebug230 @makayla_finkbeiner_ @ObeySoph▲ @YanaSidlinskiy

OpenStudy (anonymous):

so then, 325 / 0.027 = 12,037.04 pennies?

OpenStudy (anonymous):

is that right? if not how would i determine the number of pennies in the jar with this info?

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@ElisaNeedsHelp

OpenStudy (anonymous):

@Mr.Man420 @Eli_Moses @haleyu99 @GingaTheNinja<3

OpenStudy (elisaneedshelp):

@iPwnBunnies get your math butt over here

OpenStudy (eli_moses):

well, it actually depends how the pennies are placed. there can be diagonal and horizontal pennies. Each different one can take up a different space

OpenStudy (the_fizicx99):

Hm, well first I would to fine what the volume of the jar, v = π*r^2*h

OpenStudy (the_fizicx99):

find*

OpenStudy (anonymous):

ooh im not good at this o.o but @iPwnBunnies is ^-^

OpenStudy (anonymous):

well, that info isn't given so it would have to be an estimate

OpenStudy (anonymous):

penny diameter = 0.75 radius = ? 0.375 height = 0.061 base = (3.14)(0.375)^2 = 0.44 volume = (0.44)(0.061) = 0.027 glass jar diameter = 6 radius = ? 3 height = 11.5 base = (3.14)(3)^2 = 28.26 volume = (28.26)(11.5) = 324.99 | 325 all the data i calculated

OpenStudy (anonymous):

I'm just not sure if 325 / 0.027 = 12,037.04 pennies

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@RatzLover95

OpenStudy (anonymous):

is the calculation above, (below the data) correct? for the amount of pennies in ar jar? (estimate if was actual contest)

OpenStudy (the_fizicx99):

Hm, v = (3.14)*3^2*11.5 <-- this is based on your information >.> v = 3.14 * 9 * 11.5 v = 325, volume of the pennies is roughly 0.32 -- V= \pi * (0.375)^2(0.75) I tried to divide it to see how may would fit, 325/0.33 = 984.48 --> 985ish

OpenStudy (the_fizicx99):

U_U LaTeX failed ;-;

OpenStudy (anonymous):

you sure you use the diameter and not the height? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

http://prntscr.com/3c5vnp

OpenStudy (anonymous):

because the formula is (pi)(radius)^2(height)=volume

OpenStudy (the_fizicx99):

I did use the hieght

OpenStudy (anonymous):

no 0.061 is the height

OpenStudy (anonymous):

0.75 is the diameter

OpenStudy (the_fizicx99):

" Each penny is 0.75 inches in diameter and 0.061 inches"

OpenStudy (anonymous):

" diameter = 0.75 radius = ? 0.375 height = 0.061 "

OpenStudy (anonymous):

yea 0.75 is the diameter, and 0.061 is the thinkness/height

OpenStudy (the_fizicx99):

Read the problem over xD

OpenStudy (anonymous):

Each penny is 0.75 inches in diameter and 0.061 inches thick Each penny is (0.75 inches in diameter) and (0.061 inches thick)

OpenStudy (anonymous):

0.75 inches in diameter 0.061 inches thick = height

OpenStudy (anonymous):

XD wow, lol

OpenStudy (the_fizicx99):

Yup

OpenStudy (anonymous):

think i should finish this tomorrow?

OpenStudy (the_fizicx99):

Why tomorrow? :O

OpenStudy (anonymous):

i have done 3 assignments in geometry already, 2 exams and 1 activity/quiz like this

OpenStudy (anonymous):

and my dad want me to "finish"

OpenStudy (the_fizicx99):

That's up to you lol xD, My average is 11 Math assignments and Science. I view it as, why put it off if I can do it now. I am taking a break for a few mins, I just finished doing 17 assignments these past 2 days for Science. I'm done with that, I'm requesting an other course soon :P

OpenStudy (anonymous):

@Squirrels can you please help? this is really irritating... i have all the data, but i need how to use it to get the pennies in a jar

OpenStudy (the_fizicx99):

We just did it :o have fun though! Bye :> I believe Pwn is offline >.>

OpenStudy (anonymous):

@iPwnBunnies can you please help? this is really irritating... i have all the data, but i need how to use it to get the pennies in a jar